OpenStudy (anonymous):

Factor completely and then place the factors in the proper location on the grid. 36x2(the 2 is an exponent) + 60x + 25

OpenStudy (anonymous):

36x^2+60x+25 (6x+5)(6x+5)=(6x+5)^2 x=-5/6